Circular No. 7424

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000ci IN NGC 6470 T. Puckett and A. Langoussis, Mountain Town, GA, report their dis-

covery of an apparent supernova (mag 18.1) on an unfiltered CCD frame (limiting mag 20.5) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on May 12.28 UT. SN 2000ci is located at α = 17^{h}44^{m}13^{s}.21, δ = +67^{o}35′46′′.3 (equinox 2000.0), which is 1′′.7 east and 10′′.3 north of the center of NGC 6470. The new object was also present on an unfiltered CCD frame taken on May 13.12. However, SN 2000ci was not present on a frame taken on 1999 June 7.24 or on Palomar Sky Survey images taken on 1993 Aug. 17 (limiting mag ∼ 21.0) and 1987 June 2 (limiting mag ∼ 21.0).

SUPERNOVA 2000cg IN UGC 10121 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000cg (IAUC 7421), taken by M. Calkins on May 11.36 UT with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova near maximum light. The spectrum features a slightly blue continuum with broad P-Cyg profiles of Hα and Hβ. Adopting the NASA/IPAC Extragalactic Database recession velocity of 8830 km/s yields an expansion velocity for the supernova of 9000 km/s, measured via the Hα line profile.

LMC X-3 P. T. Boyd and A. P. Smale, Universities Space Research Associa-

tion and Goddard Space Flight Center, report: “RXTE/ASM observations of LMC X-3, a bright, persistent blackhole binary usually observed in the high/soft state, indicate that the source has entered a rare extended low/hard state, beginning about Apr. 10. RXTE/PCA observations on May 5.76 and 10.01 UT reveal a pure power-law spectrum with photon index 1.64 ± 0.11 and 2–10-keV flux of 8.55 × 10^{36} erg/s at 50 kpc. An additional PCA observation is planned. Radio and optical observations are encouraged so that LMC X-3 can be compared to persistent Galactic blackhole binaries that are most commonly observed in the low/hard state.”

SUPERNOVA 2000ce IN UGC 4195 CCD R magnitudes by K. Hornoch, Lelekovice, Czech Republic, with

a 0.35-m reflector: May 9.894 UT, 15.8 ± 0.1; 10.832, 16.0 ± 0.1.

2000 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green